OpenStudy (waleed_imtiaz):

cos^2(45+x) - Sin^2(45-x)= ?

OpenStudy (anonymous):

Assuming that the numbers are in degrees, you can use the phase shift identities to change a sin into a cos. But what you have up here is a unsolvable in 1 dimension, because you have two unknowns, X and ?

OpenStudy (anonymous):

Presumably wants a simplification

OpenStudy (waleed_imtiaz):

Two unknowns..... But I want to solve it...... Here see my work.... Can I write it as [cos(45+x)]^2 - [sin(45-x)]^2 ?

OpenStudy (waleed_imtiaz):

@estudier if the choices are A)0 B) 1 C)-1 D) 2

OpenStudy (waleed_imtiaz):

I think its zero..... WHat do u say ?

hartnn (hartnn):

sin (45-x)=cos (90-45+x) = cos (45+x) so answer =0

OpenStudy (waleed_imtiaz):

thanks,,,..!

hartnn (hartnn):

welcome :)